Numbering individual lines of equation array’s

Johannes Braams PTT Research Sint Paulusstraat 4 2264 XZ Leidschendam The Netherlands

         Internet: J.L.Braams@research.ptt.nl

February 9, 1994

This package defines the subeqnarray and subeqnarray* environments, which behave like the equivalent eqnarray and eqnarray* environments, except that the individual lines are numbered like 1a, 1b, 1c, etc.

To refer to these numbers an extra label command \slabel has been defined. Many of this code was taken from latex.tex and modified for this purpose.